OpenStudy (anonymous):

3g^2 - 12g = - 4 which number will complete the square? a. 12 b. -4 c. 4 d. 36

OpenStudy (anonymous):

if you factor out the 3 on the left, you'll get \(\large 3(g^2-4g)=-4\) Then to complete the square, add the square of half the coefficient of g.

OpenStudy (anonymous):

Then, consider what would have to be added to the right side to keep both sides equal.

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

Right, adding 4 inside the parentheses on the left yields \(\large 3(g^2-4g+4)=-4 + ?\)

OpenStudy (anonymous):

so 4 completes the square then?

OpenStudy (anonymous):

47

OpenStudy (anonymous):

47?

OpenStudy (anonymous):

What happens when you distribute that 3 back into those parentheses on the left?

OpenStudy (anonymous):

3g^2-12g+12

OpenStudy (anonymous):

Right, so you actually added 12 to both sides to complete the square. \(\large 3g^2−12g+12)=−4+12\). See?

OpenStudy (anonymous):

oh got it thanks!